For each of the ff. sets, determine whether 2 is an element of that set.



a. {π‘₯βˆˆβ„|π‘₯ 𝑖𝑠 π‘Žπ‘› π‘–π‘›π‘‘π‘’π‘”π‘’π‘Ÿ π‘”π‘Ÿπ‘’π‘Žπ‘‘π‘’π‘Ÿ π‘‘β„Žπ‘Žπ‘› 1}



b. {π‘₯βˆˆβ„|π‘₯ 𝑖𝑠 π‘‘β„Žπ‘’ π‘ π‘žπ‘’π‘Žπ‘Ÿπ‘’ π‘œπ‘“ π‘Žπ‘› π‘–π‘›π‘‘π‘’π‘”π‘’π‘Ÿ}



c. {2,{2}}



d. {{2},{{2}}}

1
Expert's answer
2022-03-09T15:31:05-0500

Solution (a)


The given set is {"{}" "\ud835\udc65\u2208\u211d|\ud835\udc65" 𝑖𝑠 π‘Žπ‘› π‘–π‘›π‘‘π‘’π‘”π‘’π‘Ÿ π‘”π‘Ÿπ‘’π‘Žπ‘‘π‘’π‘Ÿ π‘‘β„Žπ‘Žπ‘› 1} can be written as


"\left \{ 2, 3, 4, 5, 6, 7, 8, …\right \}"


It is clear that "2 \in \"{"{}" "\ud835\udc65\u2208\u211d|\ud835\udc65" 𝑖𝑠 π‘Žπ‘› π‘–π‘›π‘‘π‘’π‘”π‘’π‘Ÿ π‘”π‘Ÿπ‘’π‘Žπ‘‘π‘’π‘Ÿ π‘‘β„Žπ‘Žπ‘› 1}


Hence 2 is an element of that set.




Solution (b)


The given set is { "\ud835\udc65\u2208\u211d|\ud835\udc65" 𝑖𝑠 π‘‘β„Žπ‘’ π‘ π‘žπ‘’π‘Žπ‘Ÿπ‘’ π‘œπ‘“ π‘Žπ‘› π‘–π‘›π‘‘π‘’π‘”π‘’π‘Ÿ} can be written as


"\left\{ {{1^2},\,{2^2},\,{3^2},\,....} \right\}\"


"\left\{ {1,\,\,4,\,\,9,\,\,....} \right\}\"


We can see that "2 \in \left\{ {1,\,\,4,\,\,9,\,\,....} \right\}\"


Hence "2 \notin \" {π‘₯βˆˆβ„|π‘₯ 𝑖𝑠 π‘‘β„Žπ‘’ π‘ π‘žπ‘’π‘Žπ‘Ÿπ‘’ π‘œπ‘“ π‘Žπ‘› π‘–π‘›π‘‘π‘’π‘”π‘’π‘Ÿ}


Hence 2 is not an element of that set



Solution (c)


The given set is {2,{2}} can be written as


We can see that "2 \in" "\left \{ 2,\left \{2\right \} \right \}"


Hence 2 is an element of that set



Solution (d)


The given set is "\left \{ \left \{2\right \},\left \{2\right \} \right \}" can be written as


We can see that "2 \notin \
\left \{ \left \{2\right \},\left \{2\right \} \right \}"


Hence 2 is not an element of that set